if the measure of angle ABD =4x+15 and the measure of angle DBC = 2x+45 , find the degree measure of angle DBC. i just dont remember how to find x
Answer by cleomenius(959)

You need more information.
Usually in this type of problem, the angles are either complementary or supplementary. Occasionally, they may be congruent.
If they are complementary, 4x + 15 + 2x + 45 = 90.
6x + 60 = 90
6x = 30
x = 5
4x + 15 = 35 degrees.
2x + 45 = 55 degrees.
35 + 55 = 90 degrees.
=====================
If they are supplementary, 4x + 15 + 2x + 45 = 180.
6x + 60 = 180
6x = 120
x = 20
4x + 15 = 95 degrees
2x + 45 = 85 degrees
95 + 85 = 180 degrees.
==========================
If they are congruent, 4x + 15 = 2x + 45
2x = 30
x = 15 degrees.
4(15) + 15 = 75 = 2(15) + 45.
Cleomenius.